# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v2

@@ 2.3.1 Gaussian Image
To improve the reconstruction quality, several methods are subsequently proposed to leverage large models with strong capacity to learn generic scene priors from large-scale datasets for 3D scene reconstruction. Based on the 3D large reconstruction model (LRM) [10] that achieves impressive sparse-view 3D object reconstruction by learning general reconstruction priors from extensive datasets of 3D objects, GRM [79] directly maps input image pixels to a set of pixel-aligned 3D Gaussians for feed-forward 3DGS-based object reconstruction. Flash3D [80] introduces a high-quality depth predictor as prior to achieve single-view scene-level reconstruction. Concurrently, GS-LRM [81] incorporates Transformer-based LRM to formulate per-pixel Gaussian prediction as a sequence-to-sequence mapping and achieves remarkable performance across both objects and scenes. This research line is further extended by eFreeSplat [82] that leverages a large vision transformer encoder [52] as 3D priors for Gaussian image prediction, Long-LRM [83] that combines Mamba2 blocks [84] with Transformer layers to handle long sequence of input images, and several pose-free methods, including FreeSplatter [85] and AnySplat [86], that jointly predicts Gaussian images and estimates camera poses, where AnySplat leverages a pretrained VGGT [2] with geometry priors as an auxiliary supervision for the joint prediction. On the other hand, these works are constrained to reconstruct existing image observations without generative ability. To address it, LGM [87] introduces pre-trained diffusion models [88, 12, 89, 90] to generate multiview images, followed by large multiview Gaussian models to predict multiview Gaussian images. Wonderland [91] further leverages a pre-trained video diffusion model [92] to generate informative video latents from a single image for pixel-aligned 3DGS prediction.
Furthermore, to enhance the geometric quality of 3D scene reconstruction, a large number of methods incorporate geometric designs, such as epipolar and cost volumes.
@@
Epipolar-based Methods. As a pioneering epipolar-based method, PixelSplat [5] leverages an epipolar transformer to resolve the scale ambiguity issue and capture cross-view features. It then estimates a probabilistic depth distribution from the image features and predicts pixel-aligned 3D Gaussians. However, PixelSplat is effective only in regions strongly correlated with the input observations. It struggles in areas of high uncertainty, leading to blurry reconstructions that lack high-frequency details or failed reconstruction in unseen regions. LatentSplat [93] proposes to exploit a generative model to obtain high-quality reconstructions in uncertain areas. It leverages an epipolar transformer and a Gaussian sampling head to encode two-view inputs to 3D variational Gaussians and finally uses a lightweight VAE-GAN decoder [94] to generate RGB images of novel views. This approach enables high-quality binocular reconstruction of object-centric scenes with full 360° views. In addition, several methods are proposed to enable the pose-free setting. For example, GGRt [95] builds upon PixelSplat and introduces a joint learning framework for camera poses and 3D Gaussian prediction.
@@
Cost Volume-based Methods. A key limitation of PixelSplat is the inherent ambiguity and unreliability in mapping image features to depth distributions, resulting in suboptimal geometry reconstruction. To address this issue, MVSplat [96] adopts a plane-sweeping-based cost volume in 3D space to facilitate multiview Gaussian image prediction, leveraging cross-view feature similarities within the volume to provide rich geometric information for depth estimation. MVSGaussian [97] also employs a cost volume-based pipeline for pixel-aligned 3D Gaussian prediction. However, these methods heavily depend on precise multiview feature matching, which becomes particularly challenging in scenes with occlusions, low texture, or repetitive patterns. To address this issue, TranSplat [98] introduces a depth-aware deformable matching transformer to generate a depth confidence map to enhance multiview feature matching, thus improving reconstruction accuracy in areas with low texture or repetitive patterns. Similarly, DepthSplat [99] leverages robust monocular depth estimation to enhance feed-forward 3DGS reconstruction. Specifically, it combines pre-trained monocular depth features with multiview feature matching, preserving multiview depth consistency while improving robustness in these challenging scenarios. The predicted multiview depth maps are then utilized to determine the Gaussian centers, while a lightweight network estimates the remaining Gaussian parameters. Another recent work, HiSplat [100], is introduced to address the limitation of feed-forward 3DGS reconstruction in lacking hierarchical representations, which makes it difficult to simultaneously capture large-scale structures and fine texture details. After leveraging cost volume to obtain the depth and Gaussian features, HiSplat creates a large, coarse-grained Gaussian image to define the primary structure and then adds finer Gaussians around it to progressively refine and enrich the texture details. PanSplat [101] also investigates hierarchical Gaussian images for 4K panorama view synthesis. It employs a transformer-based network to build a hierarchical spherical cost volume, enabling high-resolution 3D geometry with enhanced efficiency. Besides, MVSplat360 [102] extends MVSplat to support 360° novel view synthesis for large-scale real-world scenes. For pose-free setting, Pf3plat [103] introduces a coarse-to-fine strategy to estimate the depth, confidence and camera poses and utilize them to perform 3D Gaussian prediction with the constructed multi-stereo cost volume and guidance cost volume.
@@ 2.3.2 Gaussian Volume
Gaussian volume [75] represents 3D with Gaussian voxel grids, where each voxel comprises multiple Gaussian primitives. A typical feed-forward 3DGS method using a Gaussian volume representation is LaRa [75], which aims to reduce the heavy training cost associated with 360° bounded radiance field reconstruction. As shown in Fig. 5 (b), it first builds 3D features and embedding volumes and then leverages a volume transformer to reconstruct a Gaussian volume, enabling progressively and implicitly feature matching and leading to higher quality results and faster convergence. GaussianCube [104] proposes a structured and explicit radiance representation for 3D object generation from a single image. QuickSplat [105], based on G3R [106], employs neural networks to predict voxel-level features, which are subsequently decoded into Gaussian primitives with MLPs for surface reconstruction. Besides, to enable Gaussian densification in feed-forward 3DGS, GD [107] builds upon LaRa and introduces a generative densification that exploits prior knowledge from large multiview datasets and densifies feature representations from feed-forward 3DGS. SCube [108] further advances large-scale scene reconstruction by introducing VoxSplat, a high-resolution sparse-voxel Gaussian representation generated via a hierarchical latent diffusion model conditioned on sparse posed images.
2.3.3 Gaussian Triplane
@@
Gaussian triplane refers to a hybrid 3D representation that effectively combines the high-quality representation of triplanes with the efficiency of 3DGS. It is typically constructed by triplane-based 3DGS methods, which aim to predict a triplane representation first and then leverage the latent triplane features to decode 3D Gaussians, as illustrated in Fig. 5(c). For example, Triplane-Gaussian [77] leverages several transformer-based networks pre-trained in large-scale datasets to build a Gaussian triplane, enabling high-quality single-view 3D reconstruction. AGG [109] also mixes triplane and 3D Gaussians, which first represents scene textures as triplane and then decodes 3D Gaussians from triplane-based texture features queried by 3D locations.
2.3.4 Gaussian PointMap
Gaussian pointmap refers to a hybrid 3D representation that combines pointmaps with 3D Gaussians. It is typically constructed by pointmap-based 3DGS methods, which aim to generate dense Gaussian pointmaps to enable pose-free sparse-view reconstruction and rendering. Specifically, these methods often leverage pointmaps as geometric priors, upon which 3D Gaussians are predicted, as illustrated in Fig. 5(d). With the advent of a series of feed-forward pointmap reconstruction methods [1, 4, 62], which regress dense pointmaps directly from raw unposed images, one research line of pointmap-based methods is that directly leverages the pointmap reconstruction methods to generate dense pointmaps for 3D Gaussian prediction. For example, Splatt3R [76] builds on the large-scale pretrained foundation 3D MASt3R model [4] by seamlessly integrating a Gaussian decoder, enabling pose-free feed-forward 3DGS. NoPoSplat [110] also uses MASt3R as the backbone and predicts 3D Gaussians in a canonical space without ground-truth camera poses and depth. Large spatial model [111] combines DUSt3R [1] with a Gaussian prediction head and integrates additional semantic embeddings from the input images to enable feed-forward 3D Gaussian reconstruction. SmileSplat [112] uses DUSt3R as the backbone to predict Gaussian surfels with a multi-head Gaussian regression decoder. SelfSplat [113] unifies DUSt3R-driven Gaussian prediction with self-supervised learning of depth and camera poses, enabling simultaneous prediction of geometry, pose, and Gaussian attributes. However, relying on DUSt3R and MASt3R imposes a limitation on these methods, as they inherit the constraint of pairwise inputs, restricting their scalability. PREF3R [114] builds on the pretrained reconstruction model Spann3R [62] for 3D Gaussian prediction and introduces a spatial memory network to achieve its functionality for multiview images. However, the utilization of DUSt3R, MASt3R, and Spann3R often leads to suboptimal rendering results due to imperfections in their geometry estimates.
Another research line of pointmap-based methods, FLARE [115], avoids using DUSt3R and MASt3R to obtain pointmaps, instead focusing on learning pointmaps for 3D Gaussian reconstruction. It still leverages pointmaps as the geometry representation and proposes the joint learning of camera poses, Gaussian pointmaps, enabling the high-quality feed-forward 3DGS reconstruction and rendering. Besides, LPGM [116] utilizes a pre-trained 3D diffusion model [24] to generate point clouds from a single-view input image, which are then processed by a dedicated point-to-Gaussian generator to produce the final 3D Gaussians.
@@ 2.4 Other 3D Representations
Except for the methods mentioned above, there have been several efforts dedicated to the feed-forward reconstruction with different scene representations, exploring diverse research paths. In this section, we introduce several representative and advanced methods based on mesh, occupancy, and signed distance function (SDF) representations.
@@ 2.4.1 Mesh
Meshes are compatible with various graphics pipelines and have gained significant attention in feed-forward 3D reconstruction in recent years. For example, Pixel2Mesh [117] is proposed to produce a 3D mesh from a single input image, which leverages a 2D CNN to extract image features for progressive mesh deformation. Mesh R-CNN [118] extends Mask R-CNN [119] by incorporating 3D shape inference. It introduces a voxel branch that predicts a coarse cubified mesh for each detected object, which is subsequently refined through a mesh refinement branch.
@@
Recently, one-2-3-45 [120] leverages the diffusion-based model Zero-1-to-3 [121] to produce multiview images and feeds these images to an SDF-based generalizable neural surface reconstruction module [122] for feed-forward mesh reconstruction. One-2-3-45++ [123] enhances the consistency of synthesized multiview images and utilizes a 3D diffusion-based module conditioned on multiple views to generate a textured mesh in a coarse-to-fine manner. However, they often suffer from low reconstruction quality with compromised geometry. To address this issue, Wonder3D [88] introduces a cross-domain diffusion model to generate multiview-consistent normal maps and RGB images. By leveraging these consistent outputs, it reconstructs high-quality 3D meshes through a geometry fusion process. Unique3D [124] first employs a multiview diffusion model alongside a normal diffusion model to generate multiview-consistent images and normal maps. It then introduces a fast and consistent mesh reconstruction algorithm that effectively integrates these outputs to produce high-quality 3D meshes with accurate geometry.
@@
In addition, several methods are proposed to utilize the strong capability of the large reconstruction model [10] to achieve high-quality mesh reconstruction. For example, MeshLRM [125] integrates differentiable surface extraction and rendering into a large reconstruction model, enabling the direct generation of high-fidelity 3D meshes from input images. InstantMesh [126] employs a multiview diffusion model to synthesize novel views and utilizes a transformer-based large reconstruction model to generate a high-quality 3D mesh from the multiview images. MeshFormer [11] leverages 3D voxel representations and combines 3D convolution with transformer-based LRM, leading to improved 3D mesh geometry by incorporating 3D-native designs.
@@
To generate artist-created meshes with high-quality topology, several methods draw inspiration from large language models, treating 3D meshes as sequences and introducing autoregressive transformer architectures tailored to this sequential representation. Specifically, MeshGPT [127] leverages VQVAE [128] to learn a mesh vocabulary and employs a decoder-only transformer to autoregressively generate triangle meshes in a sequential manner. To mitigate cumulative errors inherent in VQVAE-based sequence representations, MeshXL [129] introduces a neural coordinate field for sequential 3D mesh representation, enabling high-quality autoregressive mesh generation. However, these methods struggle to learn the shape distribution and the topology distribution simultaneously. To address this issue, MeshAnything [130] introduces shape-conditioned artist-created mesh generation. It leverages a pre-trained encoder [131] to extract shape features, which are then injected into the VQVAE-based sequence representation, eliminating the need to learn shape distribution and enabling the model to focus solely on topology distribution learning.
Figure 6: Typical frameworks of regression-based representation-free models. The samples are adapted from  [132] and  [7].
@@ 2.4.2 Occupancy
Occupancy [133, 42] refers to the property that describes whether a given point in a 3D space is inside or outside a surface or object. Several methods have been proposed to achieve feed-forward occupancy representation with generalization capabilities. For example, Any-Shot GIN [134] aims to model occupancy-based 3D implicit reconstruction. It begins with front-back depth estimation to generate depth maps for constructing a voxel-based representation and subsequently extracts 3D features from this volume to infer the occupancy of any 3D point in space. MCC [135] employs an encoder-decoder architecture to reconstruct an occupancy-based representation. It first encodes a compressed representation of the scene appearance and geometry and then utilizes the representation to predict occupancy probabilities and RGB colors for each 3D point. Additionally, Huang et al. introduce ZeroShape [136], a regression-based method for 3D occupancy reconstruction that achieves SOTA performance in zero-shot generalization by intermediate geometric representation and explicit reasoning.
@@ 2.4.3 SDF
Signed Distance Function (SDF) [137] is a mathematical function that represents the geometry of a shape or surface in space. For any point in 3D space (or 2D), the SDF returns the shortest distance from that point to the surface of the object. The sign of the distance indicates whether the point is inside or outside the object. Several methods have been proposed to enable feed-forward SDF representations. For example, Shap-E [24] transforms point clouds and RGBA input images into a sequence of latent vectors that serve as inputs for subsequent SDF prediction. SparseNeuS [122] initially builds a hierarchy of volumes that represent local surface details, which are then used to infer SDF-based surfaces through a progressive coarse-to-fine process. VolRecon [138] employs a view transformer to integrate features across multiple views and utilizes a ray transformer to estimate SDF values for points sampled along each ray. ReTR [139] also uses a transformer for SDF prediction. It instead introduces an occlusion transformer and a render transformer to fuse features and perform rendering. C2F2NeUS [140] incorporates multiview stereo (MVS) into SDF-based surface reconstruction by first constructing a hierarchy of geometric frustums for each view to capture local-to-global scene geometry. The features extracted from these frustums are then fused using a cross-view and cross-level fusion strategy to facilitate accurate SDF prediction. UFORecon [141] also achieves MVS-based SDF reconstruction. It introduces cross-view matching transformer to extract cross-view matching features to construct hierarchical correlation volumes, enabling impressive SDF-based surface reconstruction under camera views with limited overlaps. To improve the reconstruction quality and training efficiency, CRM [142] incorporates geometric priors into network designs based on the spatial alignment between triplanes and the six input orthographic views. Specifically, it employs a multiview diffusion model to generate six synthesized orthographic images first and then introduces a convolutional reconstruction model to map these views to triplane features, which are subsequently decoded into SDF values.
2.5 3D Representation-Free Models
Feed-forward representation-free models aim to directly feed-forward synthesize novel views without 3D representations (e.g., NeRF and 3DGS). We broadly categorize the methods into two categories: regression-based methods (Sec. 2.5.1) and generative methods (Sec. 2.5.2).
2.5.1 Regression-based Feed-Forward View Synthesis
Regression-based feed-forward methods aim to formulate the rendering process as a regression problem, learning a rendering function (typically transformer-based neural network) to predict pixel colors of novel views from sparse-view inputs directly, without relying on 3D representations like NeRF or 3DGS. The key advantage of these methods is their ability to eliminate the inductive bias inherent in 3D representations. Based on their architecture, we classify these methods into two categories: encoder-decoder models and decoder-only models.
Encoder-Decoder Models. Scene representation transformer (SRT) [132], as a representative encoder-decoder model illustrated in Fig. 6(a), leverages a transformer-based encoder to map multiview input images to latent representations first and then outputs novel-view images from a transformer-based decoder with light field rays. RUST [143] inherits an encoder-decoder architecture and enables novel view synthesis solely from RGB images, without the need for camera poses. OSRT [144] focuses on object-centric 3D scenes and incorporates a slot attention module on SRT to map the encoded latent representations to object-centric slot representations. To extend SRT to large-scale scenes, RePAST [145] integrates relative camera pose information into the attention layer of SRT. However, these methods often suffer from degraded details and suboptimal rendering quality. To address this issue, several approaches incorporate geometric information to enhance model performance. For example, GPNR [146] integrates epipolar geometry within its encoder-decoder architecture, while Du et al. [147] introduce a multiview vision transformer and epipolar line sampling to improve scene geometry. GBT [148] incorporates ray distance-based geometry reasoning into multihead attention layers of transformers in the encoder and decoder. GTA [149] introduces geometric transform attention to embed the geometrical structure of tokens into the transformer and integrates it into SRT to enhance transformer-based rendering. However, despite the improved model performance, geometrical designs often integrate additional 3D inductive biases. LVSM [7] removes the geometrical designs and leverages a transformer-based large reconstruction model with self-attention to enhance the capacity of the encoder-decoder architecture and takes posed input images and Plücker ray embeddings to regress the target view pixels. To enable 3D view synthesis without any 3D supervision—such as ground-truth 3D geometry and camera poses—RayZer [150] is proposed. It adopts the encoder-decoder architecture of LVSM [7] and introduces a large, self-supervised multiview 3D model that first learns camera parameters and latent scene representations from unposed input images, and then renders novel views.
Decoder-Only Models. To minimize 3D inductive bias introduced by latent representations in encoder-decoder architectures, LVSM [7], as illustrated in Fig. 6(b), also adopts a decoder-only design with a single-stream transformer, directly mapping input tokens to target view tokens.
@@ 2.5.1 Regression-based Feed-Forward View Synthesis
Figure 7: Representative frameworks of generative representation-free models. The samples are adapted from  [121] and  [151].
2.5.2 Generative Feed-Forward View Synthesis

### v3

@@ 2.3.1 Gaussian Map
A Gaussian map refers to a 2D-based representation of 3D Gaussians, where each 2D location encodes a 3D Gaussian. As a pioneering effort, Splatter Image [74] employs a U-Net encoder-decoder architecture [78] to predict pixel-aligned 3D Gaussians for single-view 3D object reconstruction as illustrated in Fig. 5. To improve the reconstruction quality, subsequent methods leverage generic scene priors learned from large-scale datasets. GRM [79] directly maps input image pixels to pixel-aligned 3D Gaussians for feed-forward object reconstruction, based on LRM [10] that learns general reconstruction priors from large-scale datasets of 3D objects; Flash3D [80] further introduces a high-quality depth predictor as prior to achieve single-view scene-level reconstruction. Concurrently, GS-LRM [81] incorporates Transformer-based LRM to formulate per-pixel Gaussian prediction as a sequence-to-sequence mapping, achieving remarkable performance across both objects and scenes. This line of research has been further advanced by several works yuelei here: eFreeSplat [82] leverages a large vision transformer encoder [52] as 3D priors for Gaussian image prediction; Long-LRM [83] that combines Mamba2 blocks [84] with Transformer layers to handle long sequences of input images; and FreeSplatter [85] achieves pose-free Gaussian map prediction. However, these works are limited to reconstructing existing image observations without generative capabilities. To overcome this limitation, LGM [86] introduces pre-trained diffusion models [87, 12, 88, 89] to generate multiview images, which are then used by large multiview Gaussian models for multiview Gaussian prediction. Wonderland [90] further enhances this approach by leveraging a pre-trained video diffusion model [91] to generate informative video latents from a single image for 3DGS prediction. In addition to the methods above, another line of work focuses on the geometric quality of 3D scene reconstruction by incorporating geometric designs or priors, such as epipolar constraints, cost volumes, and pre-trained reconstruction models.
@@
Epipolar-based Methods. As a pioneering epipolar-based method, PixelSplat [5] leverages an epipolar line to resolve the scale ambiguity issue and capture cross-view features, estimating a probabilistic depth distribution as 3D Gaussian positions. Although PixelSplat is effective in regions strongly correlated with the input observations, it struggles in areas of high uncertainty, leading to blurry or failed reconstructions. To address this, LatentSplat [92] combines the strengths of regression-based and generative approaches to obtain high-quality reconstructions in uncertain areas. Building further on PixelSplat, GGRt [93] introduces a joint learning framework that integrates pose estimation with a 3DGS prediction network to enable pose-free 3D Gaussian prediction.
@@
Cost Volume-based Methods. A key limitation of PixelSplat is the inherent ambiguity and unreliability in mapping image features to depth distributions, resulting in suboptimal geometry reconstruction. To address this problem, MVSplat [94] and MVSGaussian [95] adopt a plane-sweeping-based cost volume to facilitate multiview Gaussian image prediction, leveraging cross-view feature similarities to improve depth estimation. However, these methods heavily depend on precise multiview feature matching, which becomes particularly challenging in scenes with occlusions, low texture, or repetitive patterns. To address this issue, TranSplat [96] introduces a depth-aware deformable matching transformer to generate a depth confidence map to facilitate multiview feature matching. Similarly, DepthSplat [97] uses a multiview depth model that leverages pre-trained monocular depth features to enhance the feed-forward 3DGS reconstruction. This line of research is further extended by several works. HiSplat [98] introduces hierarchical representations to capture large-scale structures and fine texture details; PanSplat [99] builds a hierarchical spherical cost volume for 4K panorama view synthesis; MVSplat360 [100] adapts MVSplat to support 360° novel view synthesis for large-scale scenes; and LongSplat [101] designs an online updating framework for long-sequence image streams.
Pre-trained Reconstruction Models. With the advent of feed-forward pointmap reconstruction methods [1, 4, 62] as mentioned in Sec. 2.2, several feed-forward 3DGS methods directly use pre-trained models to generate dense pointmaps for 3D Gaussian reconstruction. For example, Splatt3R [76] builds on the large-scale pretrained foundation 3D MASt3R model [4] by integrating a Gaussian decoder, enabling pose-free feed-forward 3DGS. NoPoSplat [102] also uses MASt3R as the backbone and predicts 3D Gaussians in a canonical space without ground-truth camera poses and depth. SmileSplat [103] instead uses DUSt3R as the backbone to predict Gaussian surfels with a multi-head Gaussian regression decoder. SelfSplat [104] unifies Gaussian prediction with self-supervised learning of depth and camera poses, enabling simultaneous prediction of geometry, pose, and Gaussian attributes. However, methods relying on DUSt3R and MASt3R inherit the limitation of pairwise inputs, which restricts their scalability. Thus, PREF3R [105] adopts Spann3R [62] as the geometric prior, which introduces a spatial memory network to handle multiview images. Except for feed-forward pointmap prediction models, pre-trained 3D diffusion models [25] have also been explored as a geometric prior for 3D Gaussian prediction [106].
@@ 2.3.2 Gaussian Volume
Gaussian volume [75] represents 3D with Gaussian voxel grids, where each voxel comprises multiple Gaussian primitives. A typical feed-forward 3DGS method using a Gaussian volume representation is LaRa [75], which aims to reduce the heavy training cost associated with 360° bounded radiance field reconstruction. As shown in Fig. 5 (b), it first builds 3D features and embedding volumes and then leverages a volume transformer to reconstruct a Gaussian volume, enabling progressively and implicitly feature matching that leads to higher quality results and faster convergence. Building on the notion of structured volumetric Gaussians, GaussianCube [107] proposes a structured and explicit radiance representation for 3D object generation from a single image. In parallel, QuickSplat [108], based on G3R [109], predicts voxel-level features using neural networks, which are subsequently decoded into Gaussian primitives with MLPs for surface reconstruction. To further enrich volumetric representations, GD [110] builds upon LaRa and introduces a generative densification that exploits prior knowledge from large multiview datasets and densifies feature representations from feed-forward 3DGS. SCube [111] advances this direction toward large-scale scene reconstruction by proposing VoxSplat, a high-resolution sparse-voxel Gaussian representation generated via a hierarchical latent diffusion model conditioned on sparse posed images. Most recently, AnySplat [112] constructs volumetric Gaussians by voxelizing the Gaussian maps from variable numbers of input views, yielding a unified and efficient model capable of handling both sparse and dense views.
@@
To mitigate the high computational cost of the 3D volume, an efficient triplane structure [42, 43] is also explored for Gaussian volume prediction. It is typically constructed by predicting a triplane representation first and then leveraging the latent triplane features to decode 3D Gaussians, as illustrated in Fig. 5. For example, Triplane-Gaussian [77] leverages several transformer-based networks pre-trained in large-scale datasets to build a Gaussian triplane, enabling high-quality single-view 3D reconstruction. AGG [113] also mixes triplane and 3D Gaussians, which first represents scene textures as triplane and then decodes 3D Gaussians from triplane-based texture features queried by 3D locations.
@@ 2.4 Other 3D Representations
Except for the methods mentioned above, there have been several efforts dedicated to the feed-forward reconstruction with different 3D representations. In this section, we introduce several representative types, including methods based on the mesh, occupancy, and signed distance function (SDF).
@@ 2.4.1 Mesh
Meshes are compatible with various graphics pipelines and have gained significant attention in feed-forward 3D reconstruction in recent years. For instance, Pixel2Mesh [114] generates a 3D mesh from a single image by extracting features with a 2D CNN and progressively deforming an initial mesh; Mesh R-CNN [115] introduces a voxel branch based on Mask R-CNN [116] to predict a coarse cubified mesh followed by further refinement.
@@
This field has been further advanced by the development of diffusion models. For example, One-2-3-45 [117] builds on the diffusion-based model to generate multiview images, which are then processed by a generalizable surface reconstruction module [118] for mesh reconstruction; One-2-3-45++ [119] further utilizes a 3D diffusion-based module conditioned on multiple views to generate a textured mesh in a coarse-to-fine manner. To improve geometry consistency, Wonder3D [87] introduces a cross-domain diffusion model to generate multiview-consistent normal maps and RGB images simultaneously, which can be used for mesh reconstruction. Similarly, Unique3D [120] employs a multiview diffusion model alongside a normal diffusion model to generate multiview-consistent images and normal maps, which are fed into a fast and consistent mesh reconstruction module.
@@
In addition, several methods are proposed to take advantage of large reconstruction model [10] to achieve high-quality mesh reconstruction. For example, MeshLRM [121] integrates differentiable surface extraction and rendering into a large reconstruction model, enabling direct 3D mesh generation; InstantMesh [122] utilizes a transformer-based large reconstruction model to generate a high-quality 3D mesh from multiview images; MeshFormer [11] leverages 3D voxel representations and combines transformers with 3D convolutions to improve 3D mesh geometry.
@@
To generate artist-created meshes with high-quality topology, several methods draw inspiration from large language models, treating 3D meshes as sequences and introducing autoregressive transformer architectures tailored to this sequential representation. Specifically, MeshGPT [123] leverages VQVAE [124] to learn a mesh vocabulary and employs a decoder-only transformer to autoregressively generate triangle meshes. To mitigate cumulative errors inherent in VQVAE-based methods, MeshXL [125] further introduces a neural coordinate field for sequential 3D mesh representation. To address the problem that the above methods struggle to learn the shape and topology distributions together, MeshAnything [126] uses a pre-trained encoder [127] to inject shape features into a VQVAE-based sequence, removing the need to learn shape distribution and enabling focused topology learning.
@@ 2.4.2 Occupancy
Occupancy [128, 42] refers to the property that describes whether a given point in a 3D space is inside or outside a surface or object. Several methods have been proposed to achieve feed-forward occupancy prediction from images. For example, Any-Shot GIN [129] aims to model occupancy-based 3D implicit reconstruction. It begins with front-back depth estimation to generate depth maps for constructing a voxel-based representation, subsequently extracting 3D features from this volume to infer the occupancy of any 3D point in the space; MCC [130] encodes a compressed representation of the scene appearance and geometry, and then uses the representation to predict occupancy probability and RGB color for each 3D point; ZeroShape [131] employs intermediate geometric representation and explicit reasoning to achieve 3D occupancy regression. While there are also many works on semantic occupancy prediction for autonomous driving, we do not cover these methods here, as the focus of this survey is on 3D reconstruction and view synthesis.
@@ 2.4.3 SDF
Signed Distance Function (SDF) [132] is a mathematical function that represents the geometry of a shape or surface in space. For any point in 3D space (or 2D), the SDF returns the shortest distance from that point to the surface of the object. The sign of the distance indicates whether the point is inside or outside the object. Several methods have been proposed to enable feed-forward SDF representations. For example, SparseNeuS [118] initially builds a hierarchy of volumes that represent local surface details, which are then used to infer SDF-based surfaces through a progressive coarse-to-fine process; Shap-E [25] transforms point clouds and RGBA input images into a sequence of latent vectors for SDF prediction. Some works also incorporate multiview stereo (MVS) for SDF prediction. For instance, C2F2NeUS [133] constructs a hierarchy of geometric frustums to capture local-to-global geometry for SDF prediction; UFORecon [134] introduces a cross-view matching transformer to extract cross-view matching features to construct hierarchical correlation volumes, enabling SDF-based reconstruction with limited camera view overlaps. Without relying on MVS, CRM [135] predicts SDF by incorporating geometric priors into network designs based on the spatial alignment between triplanes and six input orthographic views generated by a multiview diffusion model. There have also been works that explore transformers for SDF prediction, including VolRecon [136] which employs a view transformer to integrate cross-view features and a ray transformer to estimate SDF values, and ReTR [137] which introduces an occlusion transformer and a render transformer to fuse features and perform rendering.
@@
Figure 6: Typical frameworks of regression-based 3D-free models. The image is adapted from  [7].
2.5 3D-Free Models
Feed-forward 3D-free models aim to directly synthesize novel views without 3D representations (e.g., NeRF and 3DGS). We broadly categorize the methods into two categories: regression-based methods (Sec. 2.5.1) and generative methods (Sec. 2.5.2).
2.5.1 Regression-based View Synthesis
Regression-based methods aim to formulate the rendering process as a regression problem, learning a rendering function (typically a transformer-based network) to predict the pixel colors of novel views from sparse-view inputs directly, eliminating the inductive bias inherent in 3D representations as shown in Fig. 6.
Scene representation transformer (SRT) [138] leverages a transformer-based encoder to map multiview input images to latent representations first and then outputs novel-view images from a transformer-based decoder with light field rays. RUST [139] inherits an encoder-decoder architecture and enables novel view synthesis solely from RGB images, without the need for camera poses. With a focus on object-centric 3D scenes, OSRT [140] incorporates a slot attention module on SRT to map the encoded latent representations to object-centric slot representations. To extend SRT to large-scale scenes, RePAST [141] integrates relative camera pose information into the attention layer of SRT. However, these methods often suffer from degraded details and suboptimal rendering quality. To address this issue, several approaches incorporate geometric information to improve model performance. For example, GPNR [142] integrates epipolar geometry within its encoder-decoder architecture, while Du et al. [143] introduce a multiview vision transformer and epipolar line sampling to improve scene geometry. GBT [144] incorporates ray distance-based geometry reasoning into multihead attention layers of transformers in the encoder and decoder. GTA [145] introduces geometric transform attention to embed the geometrical structure of tokens into the transformer and integrates it into SRT to enhance transformer-based rendering. However, despite the improved model performance, geometrical designs often integrate additional 3D inductive biases. LVSM [7] removes the geometrical designs and leverages a transformer-based large reconstruction model with self-attention to regress the target view pixels. To enable novel view synthesis without ground truth of cameras, RayZer [146] introduces a self-supervised multiview model that first learns camera parameters and latent scene representations from unposed input images, and then renders novel views.
Figure 7: Representative frameworks of generative 3D-free models. The samples are adapted from  [147] and  [148].
2.5.2 Generative View Synthesis